TRIGONOMETRIC FUNCTIONS
1.Given that sin0=4/5 and 0 lies in quadrant ll,find the following value, cos 0
2.Given that sin0=3/5 and 0 lies in quadrant ll,find the following value,sec 0
3.Given that sin=-1/2 and 0 lies in quadrant lV,find the following value,cos 0
4.Given that sin0=-1/2 and 0 lies in quadrant lV, find the following value,tan 0





















Answer by Alan3354(69443)

1.Given that sin0=4/5 and 0 lies in quadrant ll,find the following value, cos 0
cos = sqrt(1 - sin^2)
cos = 3/5
In quad II --> -3/5
---------------------
2.Given that sin0=3/5 and 0 lies in quadrant ll,find the following value,sec 0
Sec = 1/cos = -5/3
---------------------
3.Given that sin=-1/2 and 0 lies in quadrant lV,find the following value,cos 0
cos = +1/2
------------
4.Given that sin0=-1/2 and 0 lies in quadrant lV, find the following value,tan 0
tan = -sqrt(3)/3